OpenStudy (felecia):

Solve for x the following equations. -x = 6 2x - 8 = -x + 4 2x + 1/2 = 2/3 x/3 + 2 = 5 -5/x = 2

OpenStudy (kittiwitti1):

\(\color{blue}{\text{Originally Posted by}}\) @Felecia Solve for x the following equations. 1 \(\rightarrow\) -x = 6 2 \(\rightarrow\) 2x - 8 = -x + 4 3 \(\rightarrow\) 2x + 1/2 = 2/3 4 \(\rightarrow\) x/3 + 2 = 5 5 \(\rightarrow\) -5/x = 2 \(\color{blue}{\text{End of Quote}}\) 1. \(-x = 6\) is also \((-1) x = 6\) 2. Try getting the \(2x\) and the \(-x\) alone on one side. 3. Try to get the \(2x\) alone on one side. 4. Try to get the \(\Large{\frac{x}{3}}\) alone on one side. 5. If \(\Large{-\frac{5}{x}}=2\)... can you flip this around so \(x\) is alone?